# Unveiling the Invisible: Reasoning Complex Occlusions Amodally with AURA

### v1

@@ Abstract
Amodal segmentation aims to infer the complete shape of occluded objects, even when the occluded region’s appearance is unavailable. However, current amodal segmentation methods lack the capability to interact with users through text input and struggle to understand or reason about implicit and complex purposes. While methods like LISA integrate multi-modal large language models (LLMs) with segmentation for reasoning tasks, they are limited to predicting only visible object regions and face challenges in handling complex occlusion scenarios. To address these limitations, we propose a novel task named amodal reasoning segmentation, aiming to predict the complete amodal shape of occluded objects while providing answers with elaborations based on user text input. We develop a generalizable dataset generation pipeline and introduce a new dataset focusing on daily life scenarios, encompassing diverse real-world occlusions. Furthermore, we present AURA (Amodal Understanding and Reasoning Assistant), a novel model with advanced global and spatial-level designs specifically tailored to handle complex occlusions. Extensive experiments validate AURA’s effectiveness on the proposed dataset. The code, model, and dataset will be publicly released.
Figure 1: We introduce AURA, a multi-modal approach designed for reasoning the amodal segmentation mask including both visible and occlusion regions based on the user’s question. AURA can deduct the implicit purpose underneath the question and response with the textual answer along with predicted amodal masks for various objects. Best viewed in color.
@@ 1 Introduction
With the development of the Segment Anything Model (SAM) series methods [22, 53, 21], predicting the segmentation mask of visible regions has been largely solved. Amodal segmentation aims to predict the entire shape of an object, including both visible and occluded regions, which is naturally challenging since visual cues for the occluded regions are unavailable. Predicting the intact shape of occluded objects is crucial for various real-world applications, such as autonomous driving [40, 45, 13], robotics [3, 6], and image editing [54, 58, 2].
@@
To address this gap, we propose a novel task named Amodal Reasoning Segmentation (ARS). As shown in Figure 1, the ARS task aims to interact with the user through textual questions, reasoning in the complex scene, and predicting the accurate segmentation masks with textual answers. For this task, we present a highly generalizable data collection pipeline and construct a new dataset, AmodalReasonSeg, which covers daily life scenes (encompassing both indoor and outdoor environments) derived from the challenging COCOA-cls dataset [66, 11]. The AmodalReasonSeg dataset facilitates object reasoning and segmentation for visible and amodal masks across diverse scenes.
@@ 2.1 Amodal Segmentation
Amodal segmentation [25] aims to predict the complete shape of an occluded object, encompassing both visible and occluded parts. Several methods [25, 40, 11, 27, 35, 34, 29] are developed by extending the methods designed for visible segmentation or object detection [5, 16, 59, 20, 17, 18], while the relative occlusion order information is leveraged by many methods [63, 61, 38, 58]. Since shape prior knowledge is an essential approach for inferring the shape of the occluded region, several methods [52, 28, 26, 12, 10, 47, 7, 31] are designed to learn and utilize it. For instance, VRSP [52] designs a codebook to store and retrieve shape priors for amodal masks. A3D [26] learns to reconstruct 3D shape priors to guide amodal segmentation. C2F-Seg [12] works in a coarse-to-fine manner equipped with a Vector-Quantized codebook [10]. Moreover, various datasets have been proposed for diverse scenarios including daily life [66, 11], indoor [57, 9, 8, 65], autonomous driving [40, 37, 42], industrial [11], shopping [29], and gaming [15]. However, existing amodal segmentation methods and datasets are limited to using images or videos as input, focusing solely on predicting amodal masks without user interaction or reasoning about implicit user purposes. In this paper, we present the first high-quality dataset for the newly proposed amodal reasoning segmentation task and introduce a Vision-Language Model-based approach specifically designed for this task.
@@ 3.3 Occlusion Learning
Occlusion Condition Encoder is designed to improve the ability of the refined embedding $e_{r}$ predicted by the Prompt Encoder for being aware of the occlusion condition of the object(s) corresponding to the [SEG]. Specifically, the $e_{r}$ is fed into the Occlusion Condition Encoder $\mathcal{E}_{OC}$ and predicts the occlusion-aware embedding $e_{oa}$ that is distinguishable about the degree of occlusion. The Occlusion Condition Encoder $\mathcal{E}_{OC}$ additionally predicts the occlusion rate $\hat{r}$ of the object, which reflects the occlusion condition comprehensively and is supervised by $\mathcal{L}_{occ-oc}$ with the ground-truth occlusion rate $r$.
@@ 4.1 Overview of the Dataset
For benchmarking the proposed amodal reasoning segmentation (ARS) task, we present the AmodalReasonSeg dataset which is built upon the COCOA-cls dataset. The AmodalReasonSeg dataset includes 3,143 images and 35,494 high-quality question-and-answer pairs corresponding to the visible and amodal segmentation masks. The visible and amodal bounding boxes and the category information are included in our dataset. For each image, there are 11.3 pairs of questions and answers on average annotated to cover diverse potential conversations. Overall, there are 2,018 images and 22,856 pairs of questions and answers in the train set and 1,125 images and 12,638 question-and-answer pairs in the test set.
@@ 4.2 Dataset Generation
The prompt template is carefully designed to require the ChatGPT-4o to generate question-and-answer pairs that satisfy the requirements to ensure the generation of high-quality data. The annotation information of objects for this image is collected and fed to ChatGPT-4o. For each image, ten question-and-answer pairs are required to be generated. Different objects are required to be organized and contained in one answer. Besides, the relative occlusion information between different objects is required to be considered when generating the question-and-answer pairs, which makes the text description in the answer helpful for discriminating the occlusion relationship between objects. The existing annotations from the original dataset of all objects in the image are organized and taken as input for the ChatGPT-4o. Finally, three human annotators check the generated questions and verify if they are qualified. For the ones that need to be improved, human annotators will manually design appropriate questions and answers for these images. Cross-check is utilized between different human annotators to ensure the quality of generated data.
@@ 5.1 Experimental Settings
Evaluation Metrics. The gIoU and cIoU metrics are used for evaluating the segmentation performance of amodal and visible masks. The gIoU is calculated by averaging the Intersection-over-Unions (IoUs) of all images, while the cIoU is calculated by the cumulative intersection dividing the cumulative union of all masks.
@@ 5.2 Amodal Reasoning Segmentation
The amodal segmentation methods [52, 12, 57] are designed for taking an image as input and predicting amodal segmentation masks for all objects but cannot understand the user’s question, reasoning, and response with textual answers, and segmentations. Besides, MLLM-based methods [23, 50, 51, 62, 41, 64] are designed to segment the visible regions of each object by taking the image and the user’s question as input.
@@ 5.3 Ablation Study
Scale of Multi-Modal LLM Backbone. The effect of the scale of the Multi-Modal LLM backbone is evaluated as shown in Table 4. The AURA-7B, AURA-13B, and AURA-Llama2-13B denote the AURA method with different Multi-Modal LLM backbones corresponding to LLaVA-7B, LLaVA-13B, and LLaVA-Llama2-13B models. These models are different variants of the LLaVA model with different scales of parameters (7 billion and 13 billion) and different generations of Llama. All experiments apart from this ablation study are based on the AURA-7B model. We can observe from Table 4 that the AURA-7B model does not perform as well as the other two models since the fewer number of parameters limits its learning ability for fine-tuning on the proposed AmodalReasonSeg dataset. Moreover, using larger models with 13 billion parameters can bring better performance since the learning capabilities of these models are better. Finally, upgrading the Llama large language model from the first to the second generation further enhances segmentation performance by generating higher-quality answers and refined embeddings used for segmentation.
Figure 5: Visualization results of AURA. Three cases are shown from top to bottom. GT denotes the ground truth. The first two cases show the images, ground-truth amodal and visible masks, and the predicted masks of AURA. In the last case, the predicted amodal mask by AURA for the occludee, which is the occluded toilet, is shown replacing the ground-truth visible mask.
@@ 6 Conclusion
In this work, we propose a new task named Amodal Reasoning Segmentation, aiming to predict textual answers alongside visible and amodal segmentation masks in response to the user’s input question that requires understanding and complex reasoning in the image. A generalizable dataset generation pipeline is proposed, which can generate annotations with high quality efficiently. A new dataset named AmodalReasonSeg is presented for daily scenarios where occlusions are challenging for amodal segmentation. Extensive textual question-and-answer pairs are generated, and various objects are referred to in the answers. Moreover, we propose AURA to understand the user’s implicit purpose and reason for the answer, which organizes various objects coherently with accurately predicted visible and amodal segmentation masks. Extensive experiments demonstrate the effectiveness of our proposed AURA on reasoning and segmenting the amodal masks in diverse scenarios.

### v2

@@ Abstract
Amodal segmentation aims to infer the complete shape of occluded objects, even when the occluded region’s appearance is unavailable. However, current amodal segmentation methods lack the capability to interact with users through text input and struggle to understand or reason about implicit and complex purposes. While methods like LISA integrate multi-modal large language models (LLMs) with segmentation for reasoning tasks, they are limited to predicting only visible object regions and face challenges in handling complex occlusion scenarios. To address these limitations, we propose a novel task named amodal reasoning segmentation, aiming to predict the complete amodal shape of occluded objects while providing answers with elaborations based on user text input. We develop a generalizable dataset generation pipeline and introduce a new dataset focusing on daily life scenarios, encompassing diverse real-world occlusions. Furthermore, we present AURA (Amodal Understanding and Reasoning Assistant), a novel model with advanced global and spatial-level designs specifically tailored to handle complex occlusions. Extensive experiments validate AURA’s effectiveness on the proposed dataset. The code, model, and dataset are released on this page.
Figure 1: We introduce AURA, a multi-modal approach designed for reasoning the amodal segmentation mask including both visible and occlusion regions based on the user’s question. AURA can deduce the implicit purpose underneath the question and respond with the textual answer along with predicted amodal masks for various objects. Best viewed in color.
@@ 1 Introduction
With the development of the Segment Anything Model (SAM) series methods [25, 60, 24], predicting the segmentation mask of visible regions has been largely solved. Amodal segmentation aims to predict the entire shape of an object, including both visible and occluded regions, which is naturally challenging since visual cues for the occluded regions are unavailable. Predicting the intact shape of occluded objects is crucial for various real-world applications, such as autonomous driving [46, 52, 12], robotics [2, 5, 22], and image editing [61, 65].
@@
To address this gap, we propose a novel task named Amodal Reasoning Segmentation (ARS). As shown in Figure 1, the ARS task aims to interact with the user through textual questions, reasoning in the complex scene, and predicting accurate segmentation masks with textual answers. For this task, we present a highly generalizable data collection pipeline and construct a new dataset, AmodalReasonSeg, which covers daily life scenes (encompassing both indoor and outdoor environments) derived from the challenging COCOA-cls dataset [73, 10]. The AmodalReasonSeg dataset facilitates object reasoning and segmentation for visible and amodal masks across diverse scenes.
@@ 2.1 Amodal Segmentation
Amodal segmentation [28] aims to predict the complete shape of an occluded object, encompassing both visible and occluded parts. Several methods [28, 46, 10, 30, 39, 38, 32] are developed by extending the methods designed for visible segmentation or object detection [4, 18, 66, 23, 19, 20], while the relative occlusion order information is leveraged by many methods [70, 68, 44, 65]. Since shape prior knowledge is an essential approach for inferring the shape of the occluded region, several methods [59, 31, 29, 11, 9, 54, 6, 35] are designed to learn and utilize it. For instance, VRSP [59] designs a codebook to store and retrieve shape priors for amodal masks. A3D [29] learns to reconstruct 3D shape priors to guide amodal segmentation. C2F-Seg [11] works in a coarse-to-fine manner equipped with a Vector-Quantized codebook [9]. Moreover, various datasets have been proposed for diverse scenarios, including daily life [73, 10], indoor [64, 8, 7, 72], autonomous driving [46, 43, 48], industrial [10], shopping [32], and gaming [14]. However, existing amodal segmentation methods and datasets are limited to using images or videos as input, focusing solely on predicting amodal masks without user interaction or reasoning about implicit user purposes. In this paper, we present the first high-quality dataset for the newly proposed amodal reasoning segmentation task and introduce a Vision-Language Model-based approach specifically designed for this task.
@@ 3.3 Occlusion Learning
Occlusion Condition Encoder is designed to improve the ability of the refined embedding $e_{r}$ predicted by the Prompt Encoder to be aware of the occlusion condition of the object(s) corresponding to the [SEG]. Specifically, the $e_{r}$ is fed into the Occlusion Condition Encoder $\mathcal{E}_{OC}$ and predicts the occlusion-aware embedding $e_{oa}$ that is distinguishable by the degree of occlusion. The Occlusion Condition Encoder $\mathcal{E}_{OC}$ additionally predicts the occlusion rate $\hat{r}$ of the object, which reflects the occlusion condition comprehensively and is supervised by $\mathcal{L}_{occ-oc}$ with the ground-truth occlusion rate $r$. Moreover, a Linear layer and a Sigmoid layer are employed to predict the occlusion rate $\hat{r}$ based on the refined embedding $e_{r}$.
@@ 4.1 Overview of the Dataset
For benchmarking the proposed amodal reasoning segmentation (ARS) task, we present the AmodalReasonSeg dataset, which is built upon the COCOA-cls dataset. The AmodalReasonSeg dataset includes 3,143 images and 35,494 high-quality question-and-answer pairs corresponding to the visible and amodal segmentation masks. The visible and amodal bounding boxes and the category information are included in our dataset. For each image, there are 11.3 pairs of questions and answers on average annotated to cover diverse potential conversations. Overall, there are 2,018 images and 22,856 pairs of questions and answers in the train set and 1,125 images and 12,638 question-and-answer pairs in the test set.
@@ 4.2 Dataset Generation
The prompt template is carefully designed to require ChatGPT-4o to generate question-and-answer pairs that satisfy the requirements to ensure the generation of high-quality data. The annotation information of objects for this image is collected and fed to ChatGPT-4o. For each image, ten question-and-answer pairs are required to be generated. Different objects are required to be organized and contained in one answer. Besides, the relative occlusion information between different objects is required to be considered when generating the question-and-answer pairs, which makes the text description in the answer helpful for discriminating the occlusion relationship between objects. The existing annotations from the original dataset of all objects in the image are organized and taken as input for the ChatGPT-4o. Finally, three human annotators check the generated questions and verify if they are qualified. During human verification, qualified samples are retained, while flawed ones are revised with appropriate questions and answers instead of being discarded. Common issues include incorrect object references, implausible functions, or inaccurate quantities. For the ones that need to be improved, human annotators will manually design appropriate questions and answers for these images. Cross-check is utilized between different human annotators to ensure the quality of the generated data.
@@ 5.1 Experimental Settings
Evaluation Metrics. The gIoU and cIoU metrics are used for evaluating the segmentation performance of amodal and visible masks. The gIoU is calculated by averaging the Intersection-over-Unions (IoUs) of all images, while the cIoU is calculated by dividing the cumulative intersection by the cumulative union of all masks.
@@ 5.2 Amodal Reasoning Segmentation
The amodal segmentation methods [59, 11, 64] are designed for taking an image as input and predicting amodal segmentation masks for all objects, but cannot understand the user’s question, reasoning, and response with textual answers and segmentations. Besides, MLLM-based methods [26, 57, 58, 69, 47, 71] are designed to segment the visible regions of each object by taking the image and the user’s question as input.
@@ 5.3 Ablation Study
Scale of Multi-Modal LLM Backbone. The effect of the scale of the Multi-Modal LLM backbone is evaluated as shown in Table 4. The AURA-7B, AURA-13B, and AURA-Llama2-13B denote the AURA method with different Multi-Modal LLM backbones corresponding to LLaVA-7B, LLaVA-13B, and LLaVA-Llama2-13B models. These models are different variants of the LLaVA model with different scales of parameters (7 billion and 13 billion) and different generations of Llama. All experiments, apart from this ablation study, are based on the AURA-7B model. We can observe from Table 4 that the AURA-7B model does not perform as well as the other two models since the smaller number of parameters limits its learning ability for fine-tuning on the proposed AmodalReasonSeg dataset. Moreover, using larger models with 13 billion parameters can bring better performance since the learning capabilities of these models are better. Finally, upgrading the Llama large language model from the first to the second generation further enhances segmentation performance by generating higher-quality answers and refined embeddings used for segmentation.
Figure 5: Visualization results of AURA. Three cases are shown from top to bottom with questions and ground-truth (GT) answers. The first two cases show the images, ground-truth amodal and visible masks, and the predicted masks of AURA. In the last case, the predicted amodal mask by AURA for the occludee, which is the occluded toilet, is shown replacing the ground-truth visible mask.
@@ 6 Conclusion
In this work, we propose a new task named Amodal Reasoning Segmentation, aiming to predict textual answers alongside visible and amodal segmentation masks in response to the user’s input question that requires understanding and complex reasoning in the image. A generalizable dataset generation pipeline is proposed, which can generate annotations with high quality efficiently. A new dataset named AmodalReasonSeg is presented for daily scenarios where occlusions are challenging for amodal segmentation. Extensive textual question-and-answer pairs are generated, and various objects are referred to in the answers. Moreover, we propose AURA to understand the user’s implicit purpose and reason for the answer, which organizes various objects coherently with accurately predicted visible and amodal segmentation masks. Extensive experiments demonstrate the effectiveness of our proposed AURA on reasoning and segmenting the amodal masks in diverse complex scenarios. To enhance the applicability in real-world scenarios, we plan to further optimize the efficiency of our LLM-based model and enrich our dataset’s diversity and granularity. Moreover, feature compression approaches [15, 16, 17, 41, 42] can be considered to improve the inference speed of the proposed approach.
Acknowledgments
This research was supported by the Ministry of Education of Singapore Tier3 grant MOET32022-0006. This research was also supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2024 (Project Name: Global Talent for Generative AI Copyright Infringement and Copyright Theft, Project Number: RS-2024-00398413, Contribution Rate: 10%).
[…]
Supplementary Material
In the supplementary material, we introduce the following contents, including:
  • Generalization ability is demonstrated in Section 7,
  • Experimental results in Section 8,
  • Discussions in Section 9,
  • Additional Information of the Dataset in Section 10.
7 Generalization on Cloud Occluded Region Segmentation
We showcase the proposed method’s potential of generalizing to the meteorological image, which is included to qualitatively demonstrate the potential of our method to generalize across domains. The example is illustrative with no quantitative results, and is not part of the core evaluation presented in the main paper.
To be specific, cloud cover is a major challenge in remote sensing and meteorological observations, as it obscures ground-level features that are critical for weather modeling, environmental monitoring, and climate studies. Accurate retrieval of ground information under cloud cover is essential for enhancing the quality of meteorological data and improving predictions.
Ground-level observations, such as vegetation health, urban heat distribution, and soil moisture, play a vital role in meteorological models. However, cloud coverage often introduces uncertainties in these observations, affecting the accuracy of derived climate indicators. Besides, meteorological models rely on ground information for initial boundary conditions, such as vegetation cover for evapotranspiration modeling or urban features for heat island studies. Addressing cloud-induced occlusions can significantly improve the reliability of such models. Furthermore, during extreme weather events like tropical cyclones or heavy rainfall, cloud cover is prevalent. Recovering occluded ground-level features aids in post-disaster assessment and supports real-time meteorological decision-making.
It is worth noticing that the dynamic nature of cloud cover, influenced by meteorological parameters such as wind speed, humidity, and atmospheric pressure, makes it critical to develop robust methods that can adaptively infer ground information under varying cloud conditions.
We evaluate the proposed AURA method for reconstructing ground-level information obscured by cloud cover using amodal segmentation, enabling meteorologists to retrieve critical land-surface data under cloudy conditions.
As shown in Figure 6, the effectiveness of AURA is evaluated using a cloud-occluded image from the RICE dataset [33]. This dataset is derived from the Landsat 8 OLI/TIRS dataset [50], which comprises paired images captured under cloud-occluded and non-occluded conditions. The Landsat 8 OLI/TIRS dataset, developed by NASA and the United States Geological Survey (USGS), offers high-resolution multispectral and thermal imagery to support global environmental monitoring, climate change research, and sustainable resource management. For this study, we annotated ten images from the RICE dataset with ground-truth amodal masks and question-answer pairs. AURA was fine-tuned on eight images, reserving the remaining two for validation. The same model structure and parameter settings of AURA are used, as described in Section 5 of the main paper.
Figure 6 presents a qualitative result on the validation set. The results highlight AURA’s capability to segment complete shapes despite cloud occlusions, demonstrating its potential to comprehend ground-level conditions based on meteorological knowledge.
Figure 6: The qualitative result for demonstrating the generalization ability of AURA on real-world scenarios such as climate change analysis. Best viewed in color.
8 Experiments
8.1 Experimental Settings
All of the experiments are conducted on the proposed AmodalReasonSeg dataset. No extra datasets other than the proposed AmodalReasonSeg dataset are used to ensure fairness in comparison. There are no other datasets used in our experiments. The pre-trained weights of LLaVA-7B-v1-1 and SAM-Vit-H are used as the initialized weights of the multi-modal LLM as well as the vision backbone and mask decoders, respectively.
8.2 Training and Inference Efficiency
Regarding training time, LISA requires 8 hours for either visible or amodal segmentation, whereas AURA takes only 10 hours to handle both.
Under the same testing environment, LISA achieves 0.33 fps, whereas AURA achieves 0.28 fps.
8.3 Evaluation Details
Evaluation of amodal segmentation methods. The evaluation of these models is performed by computing metrics between all predicted objects and the target object, as these models are designed to predict all objects in the scene and cannot identify which specific object should be predicted.
Explanation of the evaluation of MLLM-based methods. First, MLLM methods like LISA and LLM-Seg predict one object at a time and are evaluated per prediction–target match, ignoring any unmatched targets. Because most images contain multiple objects, these methods tend to perform suboptimally. Second, for MLLM-based methods that can predict multiple objects simultaneously, such as GVSA, OMG-LLaVA, GLaMM, and PSALM, the evaluation protocol is consistent with that used for AURA.
9 Discussions
9.1 Difference between Reasoning Segmentation and Referring Segmentation
The key difference between the reasoning segmentation and the referring segmentation tasks lies in the content of the questions and answers.
In the reasoning segmentation task, questions are abstract and do not explicitly reference specific objects. The answers include both textual elaborations and segmentation masks for the objects mentioned. For example, the question, “How can I accessorize my kitchen for cooking efficiently?" requires understanding and reasoning without direct object references. The corresponding answer might be, “Utilize the bottles[SEG] on the countertop for easy access to oils or spices, and rely on the sink[SEG] for rinsing utensils and ingredients during food preparation." This answer provides both the segmentation masks (denoted as “[SEG]") and explanatory details to address the question.
In contrast, the referring segmentation task involves questions that explicitly mention the objects to be segmented. For instance, the question, “Where is the orange? Please segment it.", explicitly identifies the object of interest. The typical answer is concise, such as “It is [SEG]", and lacks additional elaborations.
9.2 Difference between Amodal Reasoning Segmentation and other Amodal Segmentation Tasks
The proposed amodal reasoning segmentation task differs from existing amodal segmentation tasks by enabling interaction with users through textual input questions and providing elaborative textual answers.
While traditional amodal segmentation tasks—spanning semantic, instance, and panoptic levels—are limited to processing only image or video inputs, the proposed task introduces the ability to accept textual questions as input, infer the implicit purpose underneath these questions, and produce textual answers alongside segmentation masks. The ability of language interaction significantly broadens the potential for real-world applications.
9.3 Applicability of Existing Reasoning Segmentation Method on the Proposed New Task
There are various existing reasoning segmentation methods proposed, like LISA. These methods were originally designed for segmenting visible masks, and can be applied to the proposed new task - Amodal Reasoning Segmentation by training with ground-truth question-answer pairs and amodal segmentation masks.
However, methods like LISA need two major improvements for this task. First, LISA can only predict visible or amodal masks, requiring an additional decoder to handle both simultaneously. Second, LISA predicts only one object, while AURA handles multiple objects in a single inference by modifying [SEG] tokens. With these changes, LISA can only be upgraded to the “Baseline” method in Table 1 of our main paper.
Then the task-specific designs, including the proposed Occlusion Condition Encoder and the Spatial Occlusion Encoder, are required to handle the occlusion problem, which is the core difference between the traditional visible segmentation task and the focused amodal segmentation task in the paper.
Figure 7: The designed prompt template used for guiding the GPT, including the core task description for the ChatGPT-4o, the guidelines for question and answer creation, and the details about the image and objects in it.
10 Dataset
10.1 Exhibition of Various Perspectives
In the created dataset AmodalReasonSeg, there are 11.3 pairs of questions and answers annotated for each image on average. To ensure the diversity of the question-and-answer pairs provided in this dataset, we specifically add a requirement in the prompt template used for guiding ChatGPT-4o to generate question-and-answer pairs from different perspectives to cover various potentials. As shown in Figures 8 and 9, two cases of the proposed AmodalReasonSeg dataset are shown to exhibit the diversity of question-and-answer pairs in this dataset.
10.2 Prompt Template for GPT
We present the designed prompt template in Figure 7. This template specifies detailed requirements and provides comprehensive information about the image and objects, ensuring the generation of high-quality questions and answers by ChatGPT-4o. Additionally, it includes an example to clarify the requirements and assist ChatGPT-4o in understanding them effectively.
10.3 Average count of question-answer pairs per image.
The dataset contains an average of 11.3 QA pairs per image. Initially, GPT-4o generated 10 QA pairs per image, which human annotators subsequently revised and occasionally supplemented during verification to enhance data quality, resulting in a final average above 10.
Figure 8: The first example of the proposed AmodalReasonSeg dataset shows various questions and answers from diverse perspectives for the same image. Best viewed in color.
Figure 9: The second example of the proposed AmodalReasonSeg dataset shows various questions and answers from diverse perspectives for the same image. Best viewed in color.